OpenStudy (anonymous):

Two jets leave Tulsa at 3:00 p.m. traveling in opposite directions. The northbound jet is flying at a rate of 850 km/h and the southbound jet is flying at a rate of 750 km/h. At what time will they be 4000 km apart?

OpenStudy (anonymous):

Their rate of separation is 850+750 = 1600 km/hour. It will take 4000/1600 = 2.5 hours = 2:30 hours for them to separate 4000 km. 3:00 + 2:30 = 5:30 pm